# The Value of Nothing111“Nowadays people know the price of everything and the value of nothing,” The Picture of Dorian Gray Wilde (1890): Multimodal Extraction of Human Values Expressed by TikTok Influencers

### v1

@@ 4 Computational Approach
Our computational approach could be placed on three axes: (i) Masked vs. Auto-regressive Language Models, (ii) Direct vs. indirect, 2-step extraction from TikTok videos (marked d versus s, repestecively), and (iii) Using pretrained LLMs vs. fine-tuned models (marked FT). In the remainder of this section we describe each axis and the way we integrate them. Figure 3 provides a high-level illustration of the architectures.
Figure 3: Direct and indirect extraction of values from TikTok movies. Fine tuning the textual models on scripts is optional. LLMs are used in a zero/few-shot manner. The MLM must be trained on a subsample in a supervised manner.
Masked vs. Auto-regressive Language Models
The classic machine learning framework depends on the availability of a labeled dataset, facilitating training in a supervised manner. A Masked Language Model (MLM) is pre-trained to predict words that are masked at random. The MLM is then used to create a contextual embedding of an input text, pushing it forward into a classification head. In this work we experiment with both roBERTa-large (Liu 2019) and DeBERTaV2-xlarge (He, Gao, and Chen 2023) with an additional classification header performing transfer learning (Devlin 2018). We use this supervised approach on textual data only – the scripts and director notes extracted from the videos by another model.
Generative (auto-regressive) Large Language Models (LLMs) have been demonstrated to perform well on a range of language processing and multimodal tasks. LLMs do not require specific training (transfer learning) as classification is done by prompting the model with the task description (Radford et al. 2019; Brown et al. 2020; Kojima et al. 2022). State-of-the-Art models such as ChatGPT.4o (OpenAI 2024) and Gemini 1.5 Pro (Google 2024) are multimodal - allowing the user to process different modalities such as text, sound and video. In this work we use Gemini 1.5 Pro with a prompt tailored for the specific tasks of script extraction and for direct and 2-step classification (see below). In the 2-step setting we also use Mistral 7B (Jiang et al. 2023).
Specifically, the prompt contains a short formal description of Schwartz’s theory of basic values, a definition for each value, along with some examples and a request to identify the values in a given TikTok (an .mp4 file is uploaded). The exact prompts are available in Appendix A.
@@
We use an LLM as a classifier in two settings – direct and indirect (2-steps). In the direct setting we prompt the model to identify and extract the values directly from the TikTok video. In the indirect approach we first prompt an LLM to extract the full script and director’s notes from a TikTok and then extract the values from the script, either in a supervised manner using one of the MLMs or in a few-shot manner, using the exact same prompt we used in the direct approach, albeit, replacing the word ‘movie’ by the word ‘script’.
An example of a generated script is presented in Table 2.
We experiment with the different approaches for three main reasons: First, previous work reported results over text in a supervised setting. Secondly, the inner-workings of (multimodal) LLMs are opaque and different prompts and tasks may result in unstable results. Finally, while LLMs are flashy and do not require supervision, it was argued that they underperform on nuanced tasks such as the detection of emotions, figurative language, implicit hate etc. (Ziems et al. 2024).
@@
[#9 IN THE WORLD]
[with one]
[fastest]
[serves]
[240KM/H]
[I]
[never]
[return]
@@
Table 2: The script generated by the video-to-script prompt (Appendix A.2) for video ID 7341895431883967746
Transfer learning and Fine-tuning
Pretrained Language Models, either masked or auto-regressive, serve as foundational models and can be used as-is, either by prompting (LLMs) or by performing transfer learning on a small annotated sample (MLMs). However, both types of language models could be further fine-tuned on in-domain data, before the classification phase. Typically [^2] Fine-tuning can also be done as part of the classifier training., fine-tuning is done in a similar way to the pretraining, that is – given a dedicated corpora, masked language models are fine-tuned by predicting randomly masked words (Zaken, Goldberg, and Ravfogel 2022), and autoregressive models are fine-tuned on next-word prediction task (Hu et al. 2021). In this work we use the roberta-large framework to fine-tune the roBERTa-large, the DeBERTaV2-xlarge, and the Mistral LLM. We report results obtained both by the basic pretrained models and the fine-tuned models.
@@
All experiments and training were executed on Google Colab Enterprise runtime environment with an NVIDIA Tesla A100 GPU (a2-highgpu-1g configuration).
5 Results, Analysis, and Discussion
The F-scores achieved by each of the models on each of the nineteen values are presented in Figure 4 and the weighted and macro F-scores are presented in Table 3. The best overall results for values and their conflicted counterparts were achieved with the non-fine-tuned RoBERTa-large (roBERTas) with weighted F-scores of 0.49/0.42/0.47 (present/contradicted/all [$+/-/all$], respectively). These levels of classification reflect the opacity of the signal, that corresponds to a multidimensional and complex human phenomenon. In fact, even human annotators required extensive training and still found the task non-trivial. At the same time, these results are competitive with previously reported state-of-the-art results (Schroter, Dementieva, and Groh 2023; Legkas et al. 2024) achieved in less challenging settings in which (i) text, rather than video, was the original medium of the input, and (ii) the number of examples available for training was an order of magnitude larger. [^3] A direct comparison is unattainable due to the very different nature of the data in terms of source, medium and style.
Looking at specific values (Figure 4(a)), best classification was achieved for Achievement (0.71), Benevolence Dependability (0.61), and Security Personal (0.58), all by the RoBERTas model. Achievement, present in more than one third of the videos, is the most frequent value in the annotated dataset, thus more occurrences are found in training and in the test set. However, the two other values occur far less than most values, hence performance cannot be attributed to prior frequency alone.
A fine-grained analysis of the results by method and value reveals interesting patterns, suggesting a number of insights. We address some of these patterns in the remainder of this section.
(a) Present manifestations (Max. F-Score: 0.71)
(b) Contradicting manifestations (Max. F-Score: 0.54)
Figure 4: F-Scores per each model and value. Values appearing less than 30 times were not included in the BERT based MLM models, thus excluded from the analysis (appearing with score 0 in the figure).
Masked Language Models Large Language Models
@@ 5 Results, Analysis, and Discussion
Weighted+ 0.49 0.44 0.30 0.33 0.29 0.33 0.40
Macro+ 0.44 0.38 0.26 0.29 0.33 0.28 0.33
Weighted- 0.42 0.41 0.33 0.30 0.17 0.21 0.22
Macro- 0.35 0.35 0.27 0.25 0.22 0.18 0.22
Weightedall 0.47 0.43 0.31 0.33 0.26 0.30 0.35
Macroall 0.39 0.36 0.25 0.26 0.29 0.24 0.29
Table 3: Weighted and Macro F-Score results for the values ($+$), the conflicted values (--) and all ($all$). $FT$: fined-tuned model; $s$ indicates extraction from scripts (2-step extraction); $d$ indicates direct extraction from video.
MLMs vs. LLMs
The best LLM results were obtained by Geminid. However, it was outperformed by both RoBERTas and DeBERTas in all metrics and over the three partitions ($\{weighted,macro\}\times\{+,-,all\}$, Table 3). Looking at single values, we observe that Gemini outperforms the MLMs in cases in which the number of occurrences of the value is minimal, hence supervised training is hampered, see results per value in Figure 4 and compare to the value distribution Figure 2. An example is Universalism Concern in the present manifestations, and Power Resources in the contradiction manifestation.
These results are consistent with the literature, reporting that strong supervised models tend to outperform current LLMs on tasks requiring nuance, typical to the Computational Social Science (CSS) research (Ziems et al. 2024).
Pretraining vs. Fine-tuning
All the fine-tuned models fell short, comparing to their pretrained counterparts. While somewhat surprising, previous studies report similar trends in various tasks and settings, e.g., (Zaken, Goldberg, and Ravfogel 2022; Kumar et al. 2022), as well as in the best performer in she value detection shared task (Schroter, Dementieva, and Groh 2023). Given the task and data at hand, we attribute this result to the combination of two main factors: (i) The relatively small dataset used for finetuning, (ii) The large number of imbalanced classes, as presented in Figure 2. Future work will explore the impact of each of these factors.
Direct vs. 2-step Extraction
The textual models (MLMs) applied in a 2-step pipeline and trained in a supervised manner, achieved better results than Geminid applied directly on the videos. However, comparing apples to apples, considering only the LLMs, we found that the direct application of Geminid on the videos significantly outperformed the application of both Geminis and Mistrals on the scripts in the 2-step pipeline. These results highlight the importance of the visual and sound elements to the TikTok platform, and thus the potential of the direct multimodal approach.
Present vs. Conflicting Manifestations
Most present values are more frequent (Figure 2) and better detected than values in conflict (Figure 4 and Table 3). It is interesting to note that while Geminid, our top performing LLM, showed decent performance on the present values (Weighted+ F-score of 0.4, compared to 0.49 by the top performing model and 0.33 by the second best LLM), it perform poorly on the conflicting values, achieving an F-score of only 0.22, compared to the 0.42 by the top performing model and only slightly better than the other LLMs. This result is surprising: while MLM results are expected to drop due to the smaller number of labeled instances available for training, we would have expected LLMs, used in a few-shot manner, to perform comparably on present and conflicting values. We have no plausible explanation for this significant drop in performance and intend to further explore it in the future.
Video to Script Conversion
Finally, the use of the multimodal Gemini 1.5Pro to convert videos to detailed textual script (master scene heading, action, dialogue) proved beneficial for the task. We find the conversion to be reasonably accurate and the level of details provide useful markers. This conversion was vital for the application of the successful MLM. Two points are worth highlighting: First, before converging on the final prompt (see Appendix A.2) we have experimented with different prompts, producing different outputs in various formats and detail. The evaluation of the quality of the script was done in a qualitative manner by two of the authors. Second, even the state-of-the-art models often fail to capture important details. For example, in a humerus nonsensical TikTok showing the influencer air-driving a car as a part of a guessing charade game, practically air driving, floating around in the room and bumping into objects, the generated script failed to reflect the fact that the influencer was actually moving in the air – the main point of the video.
Limitations
@@
This study is not without limitations. First, we sampled influencers in seven specific genres. In our dataset, we found much variability between the genres in content as well as style. We identified genres most important for children and adolescents, yet know that TikTok is extremely varied, and other genres exist. Thus, it is yet to verify that our results generalize across TikTok. Second, we applied the Schwartz personal values theory to define and identify values. This theory was tested across cultures and in hundreds of samples across the world (Sagiv and Schwartz 2022), and is considered the leading theoretical framework of values in social science. At the same time, the theory was not constructed specifically within the context of social media, and does not focus on unique aspects of this context. Specialized theories can add information for future investigation (Shifman et al. 2025). Last, in our analysis, we utilized two MLMs, and two LLMs. the use of other models may offer different results, and future studies can extend this investigation further.
@@ 6 Conclusions and Implications
We present pipeline for the extraction of personal values from TikTok, and demonstrate their utility. We show that (supervised) BERT based models outperform LLMs in value extraction, and offer a method for conversion of multimodal TikToks into text appropriate for BERT training and analysis. We also show variability in detection between different values, and between values supported and those contradicted in the TikToks.
The methods presented have important implications for research and society. TikTok emerged as a leading platform among youth (Anderson, Faverio, and Gottfried 2023; Woodward 2024), yet at the same time produces worry among parents, educators, and decision makers (Conte et al. 2024). By applying value detection, research can move on to investigate value transmission in TikTok at scale. Such research can provide society with much needed information regarding the social environment of young people. It is important to note that advertisers and influencers can make similar use of the method, in order to target youth more effectively. We follow (De Veirman, Hudders, and Nelson 2019) in call for policy makers and educators to mitigate misuse of online platforms by enabling youth to better identify patterns of influence in social media.

### v2

@@ 4 Computational Approach
We explore a number of computational approaches, differing in two fundamental ways: (i) A direct approach vs. a two-step approach, and (ii) Supervised classification vs. unsupervised classification using a LLM.
@@
We use an LLM as a classifier in two settings: direct and indirect (2-steps). In the direct setting, we extract the values directly from the TikTok mp4 video file, either using an LLM, prompting it with the task definition in a few-shots manner, or in a supervised manner, after fusing the video, audio and textual representations (MLMs vs. LLMs and Multimodality with Masked Video Autoencoder below).
The two-step approach allows us to divide and conquer: we first convert a TikTok clip to a full textual script containing master scene heading, action, and dialogue, then extract the values from the textual script.
The textual scripts are generated by a dedicated prompt. This prompt was decided after iterative prompt engineering on a representative subsample of 18 videos. Our prompt design prioritized capturing three key video elements: visual content, verbal narration, and text captions. Two of the authors independently evaluated script quality, ensuring the conversion accurately captures all three elements. The conversion was performed with Gemini 1.5 Pro. An example of a generated script is presented in Table 2. The prompt we use to convert clips to scripts is available in Appendix A.2.
Using the generated script, the extraction on the second 2-step could be done either in a supervised manner, including continuous pretraining and fine-tuning, or in a few-shot manner, using the exact same prompt used in the direct approach, albeit, replacing the word ‘movie’ with ‘script’. The two pipelines are illustrated in Figure 3.
Figure 3: Direct and indirect extraction of values from TikTok movies. continuous pre-training the textual models on scripts is optional. LLMs are used in a zero/few-shot manner. The MLM must be trained on a subsample in a supervised manner.
MLMs vs. LLMs
State-of-the-Art models such as ChatGPT-4o (OpenAI 2024) and Gemini 1.5 Pro (Google 2024) are multimodal, processing different modalities such as text, sound and video. Using them allows for a direct approach, analyzing all aspects of the TikTok simultaneously. Moreover, the LLMs allow for a few-shot extraction, with no need for supervised training. Using an LLM, either in the direct or the 2-step pipeline, the extraction is done by prompting the model with the annotation guidelines, alongside a request to identify the values in a given TikTok (either directly from the mp4 file or from the generated script). All prompts are available in Appendix A. The LLMs used were Gemini 1.5 Pro and Mistral 7B (Jiang et al. 2023).
On the other hand, the use of masked models, either MLMs like BERT or video autoencoders, allows model training, given that annotated data is available. The supervised approach is argued to better suit tasks that require nuanced understanding of the content (Ziems et al. 2024). Masked models can be used in both the direct and the two-step pipelines. In this work we experiment with roBERTa-large (Liu 2019), DeBERTa-V2-xlarge (He et al. 2023), and the Longformer (Beltagy et al. 2020).
Masked Language Models
We use pre-trained transformer models with a binary classification head. We fine-tuned the classification head with 3-fold cross-validation with early stopping to prevent overfitting. We selected the best-performing epoch for each fold based on the F1 score, and then averaged these optimal results across all folds to obtain final performance metrics (commonly referred to as ”best-epoch averaging”, recommended by (Reimers and Gurevych 2017).)
@@
NARRATOR: #9 IN THE WORLD with one fastest serves 240KM/H I never return
@@
Table 2: The script of video ID 7341895431883967746
Multimodality with Masked Video Autoencoder
VideoMAE (Tong et al. 2022), a masked autoencoder pretrained specifically on videos, is reported to achieve impressive results even over small datasets. We extract the visual embeddings using VideoMAE, audial embeddings using Wav2Vec2-Base (Baevski et al. 2020) and the textual embeddings using the SentenceTransformer (Reimers and Gurevych 2019). Fusion is done by concatenation. These fused embeddings were then fed into an MLP classification head. We refer to this fused multimodel F-VideoMAE. For more details see Appendix C.2.
Evaluation Procedure and Metrics
We evaluate the different pipeline and settings using an the standard F1-score per value and over all values (macro and weighted). We also use the recall value in a BLEU-like manner. While the different variations of the F-score allows us to evaluate the performance of each system in the detection of each specific value, as well as the overall performance. This Recall/BLEU value, inspired by the BLEU (Papineni et al. 2002) allows us to assess the performance on the movie level, measuring how many of the values conveyed in a specific TikTok were actually recovered by the algorithm, regardless of the specific mixture of values. We report and discuss the average Recall/BLEU values.
In the unsupervised settings (direct and the 2-stage approaches) the results are computed over the whole dataset. Settings involving supervised training were conducted in a 3-fold cross-validation manner, each fold uses two thirds of the data for training and a third for testing. Results are reported only for values appearing in at least 5% of the movies, in order to allow enough data in the train/test split.
@@
All experiments were executed on Google Colab Enterprise runtime environment with an NVIDIA Tesla A100 GPU (a2-highgpu-1g configuration).
Endorsed Confronted All
W M W M W M
F-VideoMAEd 0.23 0.20 0.22 0.17 0.23 0.19
Geminid 0.46 0.40 0.26 0.27 0.4 0.37
@@
[…]
Longformer2s 0.34 0.30 0.28 0.20 0.33 0.27
DeBERTa2s 0.36 0.33 0.35 0.32 0.36 0.33
Mistral2s 0.39 0.36 0.26 0.26 0.36 0.33
RoBERTa2s 0.42 0.38 0.38 0.35 0.41 0.37
Gemini2s 0.48 0.43 0.36 0.37 0.45 0.41
Table 3: Weighted (W) and Macro (M) F-Score results for the Endorsed Values , the Confronted Values and All. The direct and two-steps settings are indicated by $d$ and $2s$ (for 2-step), respectively; PTPT indicates continuous Pre-Training.
(a) Endorsed Values
(b) Confronted Values
Figure 4: F-Scores per each model and value. Values appearing less than 5% of TikToks were not included in the BERT based MLM models, thus excluded from the analysis (appearing with score 0 in the figure).
5 Results and Analysis
Overall Results
Table 3 presents the weighted and the macro F-scores achieved by each of the models. Best results overall were obtained by Gemini2s (using the 2-stage pipeline, prompting Gemini in both stages) with 0.45 (0.41) weighted (macro) F-score. Separating endorsed and confronted values, this setting achieves best results (weighted and macro) on the endorsed values, and best macro results on the confronted values. While the RoBERTa2s achieved the highest weighted F-score on the confronted values, Gemini2s is a close second, with 0.36, compared to 0.38 by RoBERTa2s. Interestingly, the direct pipeline (Geminid) tend to underperform, compared to the use of the Gemini or the RoBERTa in a 2-stage pipeline.
These results suggest that while Gemini was trained in a multimodal manner over text and video, nuanced classification better done over texts. Indeed, the 2-stage approach harnesses the multimodal capabilities of Gemini to convert videos to detailed textual scripts (master scene heading, action, dialogue) and then harness the capabilities of the LLM as a strong few-shot learner. The detailed script generated in the first stage helps mitigating the challenges demonstrated by Ziems et al. (2024).
Results per Movie (Recall/BLEU)
Focusing on the movie level, rather than on the value level, the 2-stage pipeline proved to achieve the best results (Table 4). This is consistent with the results presented above. Gemini2s demonstrates superior performance on the endorsed values and achieves the best overall results. However, when focusing specifically on confronted values at the movie level, RoBERTa2s outperforms Gemini2s. We attribute this gap to the fact that confronted values are more opaque and therefore harder to detect without explicit supervision. At the movie level, evaluation is based on recall rather than F1; for confronted values, which are more opaque, this favors models such as RoBERTa2s that exhibit higher sensitivity. This increased sensitivity, however, is accompanied by a substantially higher rate of false positives, which leads to lower precision, as evident at the value level (Table 6). However, at the movie level, where evaluation emphasizes recall rather than precision, this reduction in precision is not directly reflected. In summary, for the confronted values, at the value level this results in comparable F1 scores to Gemini2s (Table 3), while at the movie level, the higher recall of RoBERTa2s leads to improved performance.
Model Endorsed Confronted All Values
Geminid 0.51 0.2 0.36
Mistral2s 0.51 0.26 0.39
[…]
Gemini2s 0.58 0.32 0.45
DeBERTa2s 0.39 0.39 0.39
roBERTa2s 0.43 0.44 0.43
Table 4: Recall/BLEU results. The direct and two-steps pipelines are indicated by $d$ and $2s$ (2-step), respectively; PTPT indicates continuous Pre-Training.
Results per Value
The results per value are presented in the radar plots in Figure 4. The highest F-score for endorsed values (0.69) was obtained for Achievement, followed by Security Personal and Power Resources (0.54 and 0.53, respectively) while the highest score for a confronted value (0.48) was obtained for Humility, followed by Security Personal and Face (0.45, 0.41). Indeed, these are three of the five most frequent confronted values, see Figure 2.222We note that the numbers reported here result from different settings, see detailed results in Appendix B.
Endorsed vs. Confronted Values
Most of the endorsed values are more frequent in the dataset (Figure 2) and were better detected than confronted values (Figure 4 and Table 3). All models showed a performance drop between the endorsed and confronted values. The three top performing models (pipelines), Geminid, RoBERTa2S, and Gemini2s, exhibited an interesting pattern. The performance drop for Geminid (direct pipeline) was nearly twice as large as that of Gemini2s, which itself was about three times the drop observed for RoBERTa2s. The weighted F1-scores over the endorsed values were 0.48, 0.46, and 0.42 for Gemini2s, Geminid, and RoBERTa2s, respectively. Over the confronted values, the F1-scores decreased to 0.36 ($\Delta=0.12$), 0.26 ($\Delta=0.20$), and 0.38 ($\Delta=0.04$), respectively.
This drop in performance for confronted values is also evident at the movie level, where we aggregate over all values per video (Table 4). Gemini2s maintains the best overall performance (0.45), while RoBERTa2s outperforms Gemini2s on the confronted values (0.44 vs. 0.32), reflecting its higher sensitivity in detecting these opaque manifestations. The direct pipeline Geminid suffers a large drop (0.20), consistent with the pattern observed at the value level (Table 3).
It seems the manifestations of confronted values pose a challenge even for trained annotators: their detection requires an understanding of the meaning of the value, the meaning of its absence, and the way in which its absence (or confrontation) is promoted in an ironic or subversive manner. These results suggest that LLMs fall short in processing conceptual complexity and social opposition. However, these observations highlight the advantage of the 2-stage pipeline: by combining multimodal LLMs in a divide-and-conquer manner, it mitigates some of their shortcomings, achieving more robust performance both at the value level and when aggregating per movie.
Continuous Pretraining
Continuous pretraining did not improve any of the models we used. We attribute this to the combination of two main factors: (i) The relatively small dataset used for the continuous pretraining (<<1000 videos), (ii) The large number of unbalanced classes, as presented in Figure 2. Future work should explore the impact of each of these factors.
Limitations and Discussion
This study is not without limitations. First, we have sampled 91 influencers and 885 videos from TikTok. Our manual annotation process, included two independent expert annotators, and a third consolidator. This process ensured quality, yet limited the size of the dataset and the generalization beyond the TikTok platform. While this work explores the different pipelines and methods for value extraction, future work will incorporate a larger dataset, allowing for better generalization even for values infrequent in the current dataset.
Second, we have sampled influencers from seven specific genres. In our dataset, we found much variability across the genres in content and style. The genres we explored were identified as the most important for children and adolescents. TikTok, however, is extremely diverse and other genres may become prominent. Thus, future studies must verify that our results generalize across TikTok.
A third issue stems from the movie-to-script conversion. The quality of the generated scripts was established in a qualitative manner by two of the authors. However, even state-of-the-art models may fail to capture important details. For example, a humerus nonsensical TikTok showed the influencer air-driving a car in a charade game, appearing to be practically driving in the air, floating around the room and bumping into objects. The generated script failed to reflect the fact that the influencer was depicted as actually moving in the air.
@@ 5 Results and Analysis
Forth, we applied the Schwartz personal values theory to define and identify values. This theory was tested across cultures and in hundreds of samples across the world (Sagiv and Schwartz 2022), and is considered the leading theoretical framework of values in social science. At the same time, the theory was not constructed specifically within the context of social media, and does not focus on unique aspects of this context. Specialized theories can add information for future investigation (Shifman et al. 2025).
Finally, while this work explored the use of an array of strong models – new models, announced periodically, may offer a different perspective (e.g., significant improvement of the direct pipeline).
@@ 6 Conclusions and Implications
We have explored two pipelines for the extraction of personal values from TikTok – a direct approach using a SOTA multimodal LLMs and an indirect, two-steps, approach using either multimodal LLM for both stages or an LLM in the first step and a supervised MLM in the second step. We have demonstrated that the two-steps approach mitigates some of the shortcomings of SOTA LLMs – particularly when it comes to the processing of nuanced societal commentary and values. We further analyze the performance over individual values and over endorsed and confronted values.
The results reported in this paper outperform state-of-the-art results (Schroter et al. 2023; Legkas et al. 2024) achieved in less challenging settings in which (i) text, rather than video, was the original medium of the input, and (ii) the number of examples available for training was an order of magnitude larger. [^3] A direct comparison is unattainable due to the very different nature of the data in terms of source, medium and style.
The methods presented have important implications for research and society. TikTok emerged as a leading platform among youth (Anderson et al. 2023; Woodward 2024), yet at the same time produces worry among parents, educators, and decision makers (Conte et al. 2025). By applying value detection, researchers can study the prevalence transition, and effect of different values in TikTok, at scale. Such research can provide society with much needed information regarding the social environment of young people. Importantly, rather than approaching social media influence as a monolithic concern, parents can use our value identification framework to have more nuanced conversations with their children about the content they consume. In addition, the methods presented here have important implications for AI alignment, platform governance, and design choices.
It is important to note that advertisers and influencers can make similar use of the method, in order to target youth more effectively. We follow De Veirman et al. (2019) in call for policy makers and educators to mitigate misuse of online platforms by enabling youth to better identify patterns of influence in social media.
7 Ethical Considerations
Social media platforms create a semi-public sphere where influencers actively post content not just for their friends, but for a wider audience. Influencers often use their online presence as their main source of income, putting vast effort to maximize exposure and engagement (Carpiano et al. 2023; DiResta 2024). Therefore, studying influencer culture and ecosystem is not only ‘fair-practice’ under the unobtrusive observer paradigm (Kozinets 2010), but also crucial for understanding how values are being monetized by influencers targeting children. We note that eight accounts in our data feature minors, three of these are parents accounts. The legal guardians are active participants and supporters of the other five accounts.
The dataset curated for this work is publicly available and collected only via public APIs with no violation of TikTok terms of service. The annotated dataset available at [URL] contains only the TikTok IDs and the annotations, in order to allow influencers maintaining privacy should they choose to delete their content.
The societal need to understand the ecosystem that targets and influences minors balances and mitigates these ethical concerns. Moreover, we believe that only such understanding will allow parents, educators, and policy makers, to support minors – a vulnerable population – in facing potential risks resulting from the broad use of social media.